# Growth of Structural Lengthscale in Kob Andersen Binary Mixtures: Role of medium range order

### v1

@@ III.3 Mean field caging potential as structural order parameter
To characterize the static structure of glass forming systems, we introduce the structural order parameter (SOP)Manoj_prl_2021 ; Sharma2022 . The SOP quantifies the extent to which particles are confined by the mean field caging potential resulting from interactions with neighbouring particles, based on the Ramakrishnan–Yussouff free-energy functional RamakrishnanYussouff1979 . Here, we distinguish between the macroscopic SOP $\beta\Phi$ and the microscopic SOP $\beta\phi$ by this notation.
@@
where $r_{ij}$ is the separation between particles $i$ and $j$, and $\delta=0.09\sigma_{AA}$ is the width of the Gaussian smoothing kernel. This approach yields a smooth, noise-reduced local radial distribution function compared to discrete binning.
To avoid unphysical artifacts at short distances in microscopic calculations, we consistently use an approximate expression for the direct correlation function as previously detailed in references mohit_wca ; PatelSharmaBhattacharyya2023 . Specifically, the simplified form adopted is given by
@@ III.4 Structural lengthscale calculation
To quantify the structural lengthscale, we use the microscopic caging potential described before as the structural order parameter, $\beta\phi$. Its structural correlation function is defined as follows
@@
The corresponding normalized correlation function isCubuk2017 ; LandauSchuttler1999
@@ IV Results
Figure 1: Inverse of the four-point structure factor, $1/S_{4}(\mathbf{q},\tau_{\alpha})$, (Eq.3) plotted as a function of the squared wavevector, $q^{2}$, at various temperatures. Solid lines are Ornstein–Zernike (OZ) fits. (b) Excess displacement-displacement correlation, $\Gamma_{uu}(r,\Delta t)$, (Eq.7) at different temperatures. (c) Excess correlation of bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) at different temperatures. (d) Temperature dependence of three characteristic lengthscales: the dynamical length from OZ analysis of $S_{4}(\mathbf{q},\tau_{\alpha})$, $\xi_{4}$ (Eq.5); the dynamical lengthscale from $\Gamma_{uu}(r,\tau_{\alpha})$, $\xi_{D}$ (Eq.9); and the bare static lengthcale from $\Gamma_{bare}(r)$, $\xi_{bare}$ (Eq.18). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{4}$ is 4.28, $\xi_{D}$ is 0.24, & $\xi_{bare}$ is 0.71. With a decrease in temperature, both dynamical lengthscales exhibit similar growth, while the bare static lengthscale shows weaker growth. The colour coding in (b) & (c) matches that of (a).
@@
Figure 2: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu)$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu$, calculated at times scaled by the $\alpha$-relaxation time, $t/\tau_{\alpha}$, for A-type particles at different coarse-graining lengths, $L$. (a) Result at high temperature, $T=0.70$ and (b) at low temperature, $T=0.45$. The colour coding in (b) matches that of (a).
@@
Figure 3: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu(\tau_{\alpha})$, calculated at $\alpha$-relaxation time as a function of coarse-graining length, $L$, (a) for A-type particles and (b) for B-type particles at different temperatures. The colour coding in (b) matches that of (a) (c) Temperature dependence of $L_{max}$, the value of $L$ in (a) and (b) at which $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$ peaks, shows strong growth for A particles and weaker growth for B particles. Lines are power‐law fits.
@@
Figure 4: Results for 80:20 KALJ system: Top panels: Snapshots at $T=0.45$ showing structural and dynamical fields for A-type particles: (a) The bare SOP, $\beta\phi$, appears spatially noisy, lacking visible large-scale structure. (b) Coarse-graining the SOP, $\overline{\beta\phi}$, with $L_{max}$, shows the emergence of patches of high and low SOP values. (c) The particle mobility field, $\mu(\tau_{\alpha})$, calculated using Eq.19 over the $\alpha$-relaxation time shows low mobility domains that spatially mostly overlap with high SOP regions in (b), indicating a structure–dynamics link. Bottom Panels: Excess correlation functions corresponding to the fields in top panels at different temperatures. (d) The excess correlation of the bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) confirms the absence of significant growth in bare static lengthscale. (e) The excess correlation of the coarse-grained SOP, $\Gamma_{\text{CG}}(r)$, (Eq.22) shows an increasing correlation length with decreasing temperature, consistent with the emergence of extended structural patches. (f) The excess displacement–displacement correlation function, $\Gamma_{uu}(r,t_{\alpha})$, (Eq.7) exhibits growth with decreasing temperature. The colour coding in (e) and (f) matches that of (d)
@@
Figure 5: Temperature dependence of the dynamical lengthscale, $\xi_{D}$, and coarse-grained structural lengthscale, $\xi_{CG}$, in the 80:20 KALJ system. $\xi_{D}$ is obtained from $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.9); and $\xi_{CG}$ obtained from $\Gamma_{CG}(r)$, (Eq.23). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{D}$ A-type is 0.16, $\xi_{D}$ B-type is 0.15, $\xi_{CG}$ A-type is 0.91, & $\xi_{CG}$ B-type is 7.70.
@@
Figure 6: (a) Distribution of the bare SOP, $\beta\phi$, for all particles (black, circle) and for those belonging to 11A TCC clusters (red, square). The 11A particles exhibit a pronounced shift toward deeper cages. (b) Overlap function (Eq.2) for all particles (black, circle), 11A particles (red, square), and the 2000 most deeply caged particles (top 12.5%) (green, triangle). Both 11A and 2000 deeply caged particles show significantly slower dynamics than the average.
@@
Figure 7: Excess displacement–displacement correlation $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.7) at different temperatures in the 60:40 mixture. (a) A-type particles.(b) B-type particles. The colour coding in (b) matches that of (a).
@@
[…]
Figure 8: Results for 60:40 KALJ system: Top panels: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained structural order parameter, $\overline{\beta\phi}$, and mobility, $\mu(\tau_{\alpha})$, calculated at the $\alpha$-relaxation time for various coarse-graining lengths $L$, for (a) A-type particles at various temperatures and (b) B-type particles at various temperatures. (c) Temperature dependence of $L_{max}$, the value of $L$ in (a) and (b) at which $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha})$ peaks, shows strong growth for A particles and B particles. Lines are power law fits. Middle panels: Snapshot at T=0.45 for the structural fields: (d) bare SOP, $\beta\phi$, for A-type particles; (e) coarse-grained SOP, $\overline{\beta\phi}$, for A-type particles; (f) bare SOP, $\beta\phi$, for B-type particles; and (g) coarse-grained SOP, $\overline{\beta\phi}$, for B-type particles. Bare SOP appears spatially noisy, lacking visible large scale structure. Coarse-graining the SOP, with $L_{max}$, shows the emergence of patches of high and low SOP values. Bottom panels: Excess correlation functions corresponding to the fields in the middle panels at different temperatures: (h) bare correlation function, $\Gamma_{\text{bare}}(r)$ (Eq.17), for A-type particles; (i) coarse-grained correlation function, $\Gamma_{\text{CG}}(r)$ (Eq.22), for A-type particles; (j) bare correlation function, $\Gamma_{\text{bare}}(r)$ (Eq.17), for B-type particles; and (k) coarse-grained correlation function, $\Gamma_{\text{CG}}(r)$ (Eq.22), for B-type particles. While bare SOP fields show little spatial correlation due to their noisy nature, the coarse-grained fields exhibit clear growth in spatial correlations, consistent with the emergence of structural patches. The colour coding in (b),(h),(i),(j) and (k) matches that of (a).
Figure 9: Temperature dependence of the dynamical lengthscale, $\xi_{D}$, and coarse-grained structural lengthscale, $\xi_{CG}$, in the 60:40 KALJ system for the A and B types of particles. $\xi_{D}$ is obtained from $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.9); and $\xi_{CG}$ is obtained from $\Gamma_{CG}(r)$, (Eq.23). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{D}$ A-type is 1.17, $\xi_{D}$ B-type is 0.28, $\xi_{CG}$ A-type is 0.16, & $\xi_{CG}$ B-type is 0.15.
@@
There are different formalisms to calculate the dynamical lengthscale. In Fig.1(a) we plot $1/S_{4}(\mathbf{q},\tau_{\alpha})$ (Section III.1) as a function of $q^{2}$ and from this fitting using Ornstein Zernike expression (Eq.5) we obtain the dynamical correlation length $\xi_{4}$. Using the displacement-displacement correlation function, $g_{uu}(r,\tau_{\alpha})$, we calculate the excess displacement-displacement correlation function, $\Gamma_{uu}(r,\tau_{\alpha})$ (Eq.7) plotted in Fig.1(b). Integration of this excess correlation gives us another measure of the dynamic lengthscale, $\xi_{D}$ (Eq.9) Poole1998PhysicaA .
@@
Here, we study the static lengthscale as described by the recently developed structural order parameter, the depth of the mean field caging potential, $\beta\phi$  Manoj_prl_2021 ; Sharma2022 . Like the dynamical lengthscale, in Fig.1(c), we plot the excess correlation of the bare SOP, given by $\Gamma_{bare}(r)$ (Eq.17). The lengthscale $\xi_{bare}$ is then obtained from the area under the curve (Eq.18). In Fig.1(d), we plot the two dynamical lengthscales and the static lengthscale. We find that both the dynamical lengthscales grow similarly; however, the static lengthscale, almost does not grow.
@@
In order to further investigate how the structural and the dynamical lengthscales correlate, we first examine the structure-dynamics correlation. This correlation is studied using isoconfigurational ensembles. Isoconfigurational ensembles (Section III.5) were designed by Harrowell and coworkers harrowell to disentangle the influence of the initial structure from thermal noise in the dynamics. In this part of the analysis, to obtain cleaner statistics, we separately deal with the A and the B particles in the system. The results of the structure-dynamics correlation, given by the Spearman rank correlation, $C_{R}({\beta\phi},\mu)$ between the bare SOP and the mobility for the $A$ particles are given in the main text and that for the $B$ particles are given in Appendix II. As shown in Fig.2(a), $C_{R}({\beta\phi},\mu)$ between the local per particle level bare SOP with the mobility for the $A$ particles is high at shorter times, but at longer times there is a drop in the correlation. Note that we define. This drop is stronger at lower temperatures as seen in Fig.2(b). The results are similar for the $B$ particles (Fig.11 in Appendix II). This indicates that the local bare SOP has strong predictive power for short time dynamics, but its influence diminishes at longer times, particularly at low temperatures.
@@
As mentioned in Section III.1 and Section III.2, the dynamical heterogeneity lengthscale is typically evaluated at the time scale, $\tau_{\alpha}$ where the spatial correlations of particle displacements are most pronounced and dynamic heterogeneity reaches its maximum. Berthier_Biroli_PRE_2005 Thus, any static order parameter which does not show a good correlation at that timescale may not be able to predict the correct static length scale.
@@
[…]
@@
where the sum over $j$ includes all particles within a distance $L$ of the tagged particle $i$. This coarse-graining reflects the influence of a particle’s local environment, weighted exponentially by spatial proximity, and thus captures collective structural effects beyond the single-particle scale. In Fig.2(a), we plot the Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu)$ against scaled time $t/\tau_{\alpha}$ for different values of coarse-graining length. We find that at short times for higher $L$ values, the correlation is poor, and then at longer times, the correlation increases, and at even longer times, it decreases. We also find that this increase in correlation at longer times initially increases with $L$ and eventually, for higher $L$ values, the correlation drops. This implies there is an optimum value of $L$, we refer to as $L_{max}$ where the correlation is maximum. When we compare the plots in Fig.2(a) and 2(b), we find that $L_{max}$ increases with decreasing temperature. It is evident from Fig.2 that the value of $L_{max}$ also depends on the time at which the mobility is calculated.
@@
As the dynamical lengthscale is calculated at $\tau_{\alpha}$, we next analyse the correlation and the corresponding value of $L_{max}$ for the mobility at $t=\tau_{\alpha}$. In Fig.3(a) and Fig.3(b) we plot Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained SOP and the mobility calculated at $t=\tau_{\alpha}$ for the $A$ and $B$ particles, respectively, against different coarse-graining length, $L$. This plot clearly shows that at each temperature, there is an optimum value of $L$ for which the correlation is maximum. Both for A and B particles, this $L_{max}$ value increases with a decrease in temperature.
@@
However, at low temperatures, the $L_{max}$ value for the $B$ particles is less than that of the $A$ particles (Fig.3(c)), and this difference increases with a decrease in temperature. Note that although we study the A and the B particles separately, while coarse-graining the SOP, if the neighbouring particle is of a different species, its value is included in the coarse-graining process. Thus, the lower value of $L_{max}$ for the $B$ particles is not due to a lower concentration of the $B$ particles. This just shows that, compared to that for the A particles, fewer particles contribute to the long time dynamics of the $B$ particles.
@@
In Fig.4(a-c), we show the snapshots of the single particle bare SOP, coarse-grained SOP and also the mobility at $t=\tau_{\alpha}$ for the A particles at $T=0.45$. We find that the snapshot of the single particle bare SOP (Fig.4(a)) is highly fluctuating data and has no correlation with the mobility at long times (Fig.4(c)). However, when we plot SOP, which is coarse-grained over $L_{max}=3.56$ we find regions of high and low values of SOP (Fig.4(b)) which visually correlates with the regions of high and low values of mobility (Fig.4(c)) It is this correlation between the coarse-grained SOP and the mobility that is quantified in terms of Spearman rank correlation values given in Fig.3(a). From the above plots, a key observation is that once we introduce a coarse-graining length $L_{max}$, the ability of SOP to predict which particles will move significantly over long times improves. Physically, this indicates that the long time rearrangements depend not just on the SOP value of a single particle but also on the SOP values of its surrounding particles, giving rise to the medium range order. The lengthscale of this medium range order also increases with a decrease in temperature, a concept which is in tune with the growing cooperatively rearranging regions.
@@
Thus, to extract the correct static lengthscale which is commensurate with the dynamical lengthscale we reassign the SOP value of each particle with its coarse-grained value and calculate the excess correlation between the coarse-grained SOP as,

### v2

@@ III.3 Mean field caging potential as structural order parameter
To characterise the static structure of glass forming systems, we introduce the structural order parameter (SOP)Manoj_prl_2021 ; Sharma2022 . The SOP quantifies the extent to which particles are confined by the mean field caging potential resulting from interactions with neighbouring particles, based on the Ramakrishnan–Yussouff free-energy functional RamakrishnanYussouff1979 . Here, we distinguish between the macroscopic SOP $\beta\Phi$ and the microscopic SOP $\beta\phi$ by this notation.
@@
where $r_{ij}$ is the separation between particles $i$ and $j$, and $\delta=0.09\sigma_{AA}$ is the width of the Gaussian smoothing kernel. This approach yields a smooth, noise-reduced local radial distribution function compared to discrete binning. The correlation between structure and dynamics is best captured when the upper limit of the $r$-integration is set at the position of the first minimum of $g_{uv}(r)$, computed using Eq. 11. This choice ensures that only the first-nearest neighbours contribute significantly to the SOP. In addition, we impose a lower cutoff on the $r$-integration. This lower bound is selected as the distance at which the average radial distribution function first attains a non-zero valueSharma2022 .
To avoid unphysical artefacts at short distances in microscopic calculations, we consistently use an approximate expression for the direct correlation function as previously detailed in references mohit_wca ; PatelSharmaBhattacharyya2023 . Specifically, the simplified form adopted is given by
@@
Note that while writing the macroscopic caging potential, Eq.10 and its microscopic counterpart, Eq.13, we have made a mean field approximation. This implies that we have assumed the background of the tagged particle to be frozen, ignoring its dynamics. Even under this mean field approximation using local rdf, given by Eq.14, it is possible to capture the heterogeneity in the local structure of each particle and thus the distribution of the SOP as shown in Appendix I. We find that with decreases in temperature, the environment around each particle becomes more structured, and the distribution shifts to deeper caging potentials.
@@ III.4 Structural lengthscale calculation
To quantify the structural lengthscale, we use the microscopic caging potential described before as the structural order parameter, $\beta\phi$ and follow the methodology used for obtaining the dynamical lengthscale from the displacement-displacement correlation function. The excess correlation function of the structural order parameter is defined as follows
@@
The corresponding normalized correlation function is given by, Cubuk2017 ; LandauSchuttler1999
@@
For this calculation, we take 16,000 particle system. The averaging is done over a run length of 200,000 time steps for 12 configurations.
@@ IV Results
Figure 1: (a) Inverse of the four-point structure factor, $1/S_{4}(\mathbf{q},\tau_{\alpha})$, (Eq.3) plotted as a function of the squared wavevector, $q^{2}$, at various temperatures. Solid lines are Ornstein–Zernike (OZ) fits. (b) Temperature dependence of three characteristic lengthscales: the dynamical lengthscale, $\xi_{4}$ (Eq.5); the dynamical lengthscale, $\xi_{D}$ (Eq.9) for A particles; and the bare static lengthcale, $\xi_{bare}$ (Eq.18) for A particles. The scaling factor for $\xi_{4}$ is 2.21, $\xi_{D}$ is 0.31, & $\xi_{bare}$ is 0.28. We also plot the PTS lengthscale reported by Hocky et al. Hocky2012 scaled by 1.63, the PTS and finite size scaling lengthscale reported by Chakrabarty et al. Chakrabarty2017 scaled by 1.60 and the static lengthscale obtained by Zhang and Kob for the angular power spectra Zhang2020 scaled by 3.52. With a decrease in temperature, both dynamical lengthscales exhibit similar growth, while the static lengthscales show weaker growth.
@@
Figure 2: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu)$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu$, calculated at times scaled by the $\alpha$-relaxation time, $t/\tau_{\alpha}$, for A-type particles at different coarse-graining lengths, $L$. (a) Result at high temperature, $T=0.70$ and (b) at low temperature, $T=0.47$. The colour coding in (b) matches that of (a).
@@
Figure 3: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu(\tau_{\alpha})$, calculated at $\alpha$-relaxation time as a function of coarse-graining length, $L$, (a) for A-type particles and (b) for B-type particles at different temperatures. The colour coding in (b) matches that of (a) (c) Temperature dependence of the average value of $L_{max}$, shows stronger growth for A particles and weaker growth for B particles. The error bars show the variation of the $L_{max}$ with configuration. Lines are a guide to the eye.
@@
Figure 4: Results for 80:20 KALJ system: Top panels: Snapshots at $T=0.45$ showing structural and dynamical fields for A-type particles: (a) The bare SOP, $\beta\phi$, appears spatially noisy, lacking visible large-scale structure. (b) Coarse-graining the SOP, $\overline{\beta\phi}$, with $L_{max}$, shows the emergence of patches of high and low SOP values. (c) The particle mobility field, $\mu(\tau_{\alpha})$, calculated using Eq.19 over the $\alpha$-relaxation time shows low mobility domains that spatially mostly overlap with high SOP regions in (b), indicating a structure–dynamics link. Bottom Panels: Normalized excess correlation functions corresponding to the fields in top panels at different temperatures. (d) The normalised excess correlation of the bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) confirms the absence of significant growth in bare static lengthscale. (e) The normalised excess correlation of the coarse-grained SOP, $\Gamma_{\text{CG}}(r)$, (Eq.22) shows an increasing correlation length with decreasing temperature, consistent with the emergence of extended structural patches. (f) The normalized excess displacement–displacement correlation function, averaged over the configurations $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.8) exhibits growth with decreasing temperature. The colour coding in (e) and (f) matches that of (d)
@@
Figure 5: Temperature dependence of the dynamical lengthscale, $\xi_{D}$, and coarse-grained structural lengthscale, $\xi_{CG}$, in the 80:20 KALJ system. $\xi_{D}$ is obtained from $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.9); and $\xi_{CG}$ obtained from $\Gamma_{CG}(r)$, (Eq.23). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{D}$ A-type is 0.31, $\xi_{D}$ B-type is 0.22, $\xi_{CG}$ A-type is 1.00, & $\xi_{CG}$ B-type is 6.489. The dotted line is a power law fit with $\xi=\xi_{o}[(T-T_{VFT})/T_{VFT}]^{-2/3}$, where $T_{VFT}=0.319$, same as that obtained for the relaxation time. Inset: We plot $ln(\tau_{\alpha})$ against $\xi^{3/2}$ for $\xi_{4}$, $\xi_{D}$ and $\xi_{CG}$ for A particles. The linear fit confirms a power law relationship between the dynamics and the lengthscales.
@@
Figure 6: (a) Distribution of the bare SOP, $\beta\phi$, for all particles (black, circle) and for those belonging to 11A TCC clusters (red, square). The 11A particles exhibit a pronounced shift toward deeper cages. The 11A structure, which is a bicapped square antiprism, is also shown. (b) Overlap function (Eq.2) for all particles (black, circle), 11A particles (red, square), and the 2000 most deeply caged particles (top 12.5%) (green, triangle). Both 11A and 2000 deeply caged particles show significantly slower dynamics than the average.
@@
Figure 7: Normalized excess displacement–displacement correlation $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.8) at different temperatures in the 60:40 mixture. (a) A-type particles.(b) B-type particles. The colour coding in (b) matches that of (a).
@@
Figure 8: Results for 60:40 KALJ system: Top panels: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained structural order parameter, $\overline{\beta\phi}$, and mobility, $\mu(\tau_{\alpha})$, calculated at the $\alpha$-relaxation time for various coarse-graining lengths $L$, for (a) A-type particles at various temperatures and (b) B-type particles at various temperatures. (c) Temperature dependence of $L_{max}$, the value of $L$ in (a) and (b) at which $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha})$ peaks, shows strong growth for A particles and B particles. Lines are a guide to the eye.Middle panels: Snapshot at T=0.45 for the structural fields: (d) bare SOP, $\beta\phi$, for A-type particles; (e) coarse-grained SOP, $\overline{\beta\phi}$, for A-type particles; (f) bare SOP, $\beta\phi$, for B-type particles; and (g) coarse-grained SOP, $\overline{\beta\phi}$, for B-type particles. Bare SOP appears spatially noisy, lacking visible large scale structure. Coarse-graining the SOP, with $L_{max}$, shows the emergence of patches of high and low SOP values. Bottom panels: Excess correlation functions corresponding to the fields in the middle panels at different temperatures: (h) bare correlation function, $\Gamma_{\text{bare}}(r)$ (Eq.17), for A-type particles; (i) coarse-grained correlation function, averaged over the configurations, $\Gamma_{\text{CG}}(r)$ (Eq.22), for A-type particles; (j) bare correlation function, $\Gamma_{\text{bare}}(r)$ (Eq.17), for B-type particles; and (k) coarse-grained correlation function, averaged over the configurations, $\Gamma_{\text{CG}}(r)$ (Eq.22), for B-type particles. While bare SOP fields show little spatial correlation due to their noisy nature, the coarse-grained fields exhibit clear growth in spatial correlations, consistent with the emergence of structural patches. The colour coding in (b),(h),(i),(j), and (k) matches that of (a).
Figure 9: Temperature dependence of the dynamical lengthscale, $\xi_{D}$, and coarse-grained structural lengthscale, $\xi_{CG}$, in the 60:40 KALJ system for the A and B types of particles. $\xi_{D}$ is obtained from $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.9); and $\xi_{CG}$ is obtained from $\Gamma_{CG}(r)$, (Eq.23). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{D}$ A-type is 0.30, $\xi_{D}$ B-type is 0.25, $\xi_{CG}$ A-type is 1.24, & $\xi_{CG}$ B-type is 2.38. The dotted line is a power law fit with $\xi=\xi_{o}[(T-T_{VFT})/T_{VFT}]^{-2/3}$, where $T_{VFT}=0.312$, same as that obtained from the relaxation time. Inset: We plot $ln(\tau_{\alpha})$ against $\xi^{3/2}$ for $\xi_{4}$, $\xi_{D}$ and $\xi_{CG}$ for A and B particles. The linear fit confirms a power law relationship between the dynamics and the lengthscales.
@@
There are different formalisms to calculate the dynamical lengthscale. In Fig.1(a) we plot $1/S_{4}(\mathbf{q},\tau_{\alpha})$ (Section III.1) as a function of $q^{2}$ and from this fitting using Ornstein Zernike expression (Eq.5) we obtain the dynamical correlation length $\xi_{4}$. In Appendix II, we plot the $S_{4}({q},\tau_{\alpha})/S_{4}({q=0},\tau_{\alpha})$ as a function of $q\xi_{4}$ which shows a master plot following the OZ relationship. As we know from previous studies, this data collapse predicts the reliability of the OZ fitting Szamel_s4qt ; Tah_dis_dis . Using the displacement-displacement correlation function, $g_{uu}(r,\tau_{\alpha})$, we calculate the excess displacement-displacement correlation function, $\Gamma_{uu}(r,\tau_{\alpha})$ (Eq.8) for the A particles, plotted later in Fig.4(f). Integration of this excess correlation gives us another measure of the dynamic lengthscale, $\xi_{D}$ (Eq.9) Poole1998PhysicaA .
@@
Here, we study the static lengthscale as described by the recently developed structural order parameter, the depth of the mean field caging potential, $\beta\phi$  Manoj_prl_2021 ; Sharma2022 . To be consistent with our methodology, similar to the way the dynamical lengthscale is obtained from the normalised excess displacement-displacement correlation, we use the normalised excess correlation of the SOP to compute the static lengthscale as given by Eq.17. The lengthscale $\xi_{bare}$ for the A particles is then obtained from the area under the curve (Eq.18) of the normalised excess correlation of the bare SOP for the A particles shown later in Fig.4(d). In Fig.1(b), we plot the two dynamical lengthscales and the static lengthscale. We find that both the dynamical lengthscales grow similarly; however, the static lengthscale, grows much slowly. In the same figure for comparison, we also plot the PTS lengthscale reported by Hocky et al. Hocky2012 , the PTS and finite size scaling lengthscale reported by Chakrabarty et al. Chakrabarty2017 , and the static lengthscale obtained by Zhang and Kob for the angular power spectra Zhang2020 . We find that in the temperature regime that we have studied, our static lengthscale, the PTS lengthscale, and that associated with the angular power spectra grow in a similar fashion, more slowly than the dynamic lengthscales.
@@
In order to further investigate how the structural and the dynamical lengthscales correlate, we first examine the structure-dynamics correlation. This correlation is studied using isoconfigurational ensembles. Isoconfigurational ensembles (Section III.5) were designed by Harrowell and coworkers harrowell to disentangle the influence of the initial structure from thermal noise in the dynamics. In this part of the analysis, to obtain cleaner statistics, we separately deal with the A and the B particles in the system. The results of the structure-dynamics correlation, given by the Spearman rank correlation, $C_{R}({\beta\phi},\mu)$ between the bare SOP and the mobility for the $A$ particles are given in the main text, and those for the $B$ particles are given in Appendix IV. As shown in Fig.2(a), $C_{R}({\beta\phi},\mu)$ between the local per particle level bare SOP with the mobility for the $A$ particles is high at shorter times, but at longer times, there is a drop in the correlation. This drop is stronger at lower temperatures as seen in Fig.2(b). The results are similar for the $B$ particles (Fig.14 in Appendix IV). This indicates that the local bare SOP has strong predictive power for short time dynamics, but its influence diminishes at longer times, particularly at low temperatures.
@@
As mentioned in Section III.1 and Section III.2, the dynamical heterogeneity lengthscale is typically evaluated at the time scale, $\tau_{\alpha}$ where the spatial correlations of particle displacements are most pronounced and dynamic heterogeneity reaches its maximum. Berthier_Biroli_PRE_2005 Thus, any static order parameter that does not show a good correlation at that timescale may not be able to predict the correct static lengthscale.
@@
$\overline{\beta\phi}_{i}(L)=\frac{\displaystyle\sum_{j\neq i}\beta\phi(\mathbf{r}_{j})\,\exp\bigl[-|\mathbf{r}_{i}-\mathbf{r}_{j}|/L\bigr]\;+\;\beta\phi(\mathbf{r}_{i})}{\displaystyle\sum_{j\neq i}\exp\bigl[-|\mathbf{r}_{i}-\mathbf{r}_{j}|/L\bigr]\;+\;1}\,$ (21)
@@
where the sum over $j$ includes all particles within a distance $L$ of the tagged particle $i$. This coarse-graining reflects the influence of a particle’s local environment, weighted exponentially by spatial proximity, and thus captures collective structural effects beyond the single-particle scale. In Fig.2(a), we plot the configuration averaged Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu)$ against scaled time $t/\tau_{\alpha}$ for different values of coarse-graining length. We find that at short times for higher $L$ values, the correlation is poor, and then at longer times, the correlation increases, and at even longer times, it decreases. We also find that this increase in correlation at longer times initially increases with $L$ and eventually, for higher $L$ values, the correlation drops. This implies there is an optimum value of $L$, we refer to as $L_{max}$ where the correlation is maximum. When we compare the plots in Fig.2(a) and 2(b), we find that $L_{max}$ increases with decreasing temperature. It is evident from Fig.2 that the value of $L_{max}$ also depends on the time at which the mobility is calculated.
@@
As the dynamical lengthscale is calculated at $\tau_{\alpha}$, we next analyse the correlation and the corresponding value of $L_{max}$ for the mobility at $t=\tau_{\alpha}$. In Fig.3(a) and Fig.3(b) we plot the configuration averaged Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained SOP and the mobility calculated at $t=\tau_{\alpha}$ for the $A$ and $B$ particles, respectively, against different coarse-graining length, $L$. This plot clearly shows that at each temperature, there is an optimum value of $L$ for which the correlation is maximum. For the estimation of the value of $L_{max}$ instead of using the configuration averaged Spearman Rank correlation we plot the $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$ for every configuration and obtain the peak position which is the corresponding $L_{max}$ value for that configuration. We find that there is a variation of the $L_{max}$ value with configuration. The average of these $L_{max}$ values is plotted in Fig.3(c), and the spread of the data is represented by the error bars. Both for A and B particles, this average $L_{max}$ value increases with a decrease in temperature.
@@
However, at low temperatures, the $L_{max}$ value for the $B$ particles grows less than that of the $A$ particles (Fig.3(c)), and this difference increases with a decrease in temperature. Note that although we study the A and the B particles separately, while coarse-graining the SOP, if the neighbouring particle is of a different species, its value is included in the coarse-graining process. Thus, the lower value of $L_{max}$ for the $B$ particles is not due to a lower concentration of the $B$ particles. This just shows that, compared to that for the A particles, fewer particles contribute to the long time dynamics of the $B$ particles.
@@
In Fig.4(a-c), we show the snapshots of the single particle bare SOP, coarse-grained SOP, and also the mobility at $t=\tau_{\alpha}$ for the A particles at $T=0.45$. We find that the snapshot of the single particle bare SOP (Fig.4(a)) is highly fluctuating data and has no correlation with the mobility at long times (Fig.4(c)). However, when we plot SOP, which is coarse-grained over $L_{max}=3.56$ we find regions of high and low values of SOP (Fig.4(b)) which visually correlates with the regions of high and low values of mobility (Fig.4(c)) It is this correlation between the coarse-grained SOP and the mobility that is quantified in terms of Spearman rank correlation values given in Fig.3(a). From the above plots, a key observation is that once we introduce a coarse-graining length $L_{max}$, the ability of SOP to predict which particles will move significantly over long times improves. Physically, this indicates that the long time rearrangements depend not just on the SOP value of a single particle but also on the SOP values of its surrounding particles, giving rise to the medium range order. The lengthscale of this medium range order also increases with a decrease in temperature, a concept that is in tune with the growing cooperatively rearranging regions.
@@
Thus, to extract the correct static lengthscale which is commensurate with the dynamical lengthscale we reassign the SOP value of each particle with its coarse-grained value and calculate the normalised excess correlation between the coarse-grained SOP as,